INFRARED ABSORPTIONS OF THE H_{2}O—H_{2} COMPLEX TRAPPED IN SOLID NEON

DANIEL FORNEY, \underline{MARILYN E. JACOX}, AND WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

The infrared spectrum of a mixture of H_{2}O and H_{2} with a large excess of neon which has been deposited at 4.2 K includes several new absorptions which can be assigned to the H_{2}O—H_{2} van der Waals complex. The H–H stretching vibration of the H_{2} moiety is detectable even for very dilute concentrations of the complex. Small shifts in the position of the OH-stretching absorptions associated with vibrotational transitions of the H_{2}O moiety also are observed. The intensities of absorptions arising from nonrotating H_{2}O are considerably enhanced for the complex. Studies of samples enriched in D_{2}O and in D_{2} give similar results. Since HD, unlike H_{2} and D_{2}, is readily thermally depopulated to its J = 0 rotational state, this observation is consistent with the stabilization principally of the complex between H_{2}O and H_{2} or D_{2} in its J = 1 rotational state.